# Collective Variable Free Transition Path Sampling with Generative Flow Network

### v2

@@ Abstract
Understanding transition paths between meta-stable states in molecular systems is fundamental for material design and drug discovery. However, sampling these paths via molecular dynamics simulations is computationally prohibitive due to the high-energy barriers between the meta-stable states. Recent machine learning approaches are often restricted to simple systems or rely on collective variables (CVs) extracted from expensive domain knowledge. In this work, we propose to leverage generative flow networks (GFlowNets) to sample transition paths without relying on CVs. We reformulate the problem as amortized energy-based sampling over molecular trajectories and train a bias potential by minimizing the squared log-ratio between the target distribution and the generator, derived from the flow matching objective of GFlowNets. Our evaluation on three proteins (Alanine Dipeptide, Polyproline, and Chignolin) demonstrates that our approach, called TPS-GFN, generates more realistic and diverse transition paths than the previous CV-free machine learning approach.
@@ 1 Introduction
Project page: https://anonymous.4open.science/w/tps_gflow/.
@@
In material design and drug discovery, understanding the mechanisms and kinetics of transitions between meta-stable states of molecular systems, such as protein folding and chemical reactions [1, 2, 3, 4], is crucial. A comprehensive study of them requires sampling diverse transition paths [5, 6], which provide insights into mechanisms and energy landscapes. However, sampling transition paths through unbiased molecular dynamics (MD) simulations is challenging due to the high-energy barriers of intermediate states, which cause an exponential decay in transition probability [7]. Thus, a brute-force search through unbiased MD simulations is impractical within a realistic computational budget.
@@
To address this problem, researchers have developed bias potential enhanced sampling (BPES) methods such as umbrella sampling [8], meta-dynamics [9], and adaptive biasing force [10, ABF] methods. BPES methods rely on auxiliary bias potentials (or forces) to facilitate transitions across high-energy barriers. These methods design bias potentials using collective variables (CVs), functions of atomic coordinates that capture the slow modes of the transition. While effective for some systems, the reliance on expensive domain knowledge limits the applicability of BPES methods to systems where CVs are less understood.
@@
Recently, machine learning has emerged as a promising paradigm for designing bias potentials (or forces) without relying on CVs [11, 12, 13, 14]. For instance, Das et al. [11] and Lelièvre et al. [12] used reinforcement learning to sample transition paths from biased molecular dynamics with a parametrized force that approximates the transition path distribution. However, these works are limited to two-dimensional systems. Holdijk et al. [14] trained the bias potential by minimizing KL divergence on the trajectories collected by itself, i.e., on-policy training. Still, this approach suffers from mode-seeking behavior, and the on-policy training does not reuse the collected trajectories, leading to inefficiency in sample complexity.
@@
Contribution. In this work, we propose TPS-GFN, generative flow networks [15, GFlowNets] (GFN) for transition path sampling (TPS), which trains the bias potential without requiring collective variables (CVs). To this end, we formulate the problem as the energy-based sampling over molecular trajectories between the given meta-stable states. GFlowNet allows off-policy training and avoids collapsing to a particular mode of distribution over transition paths.
@@
In particular, our method trains the bias potential by minimizing the squared log-ratio between the target distribution and the generator, which we derive from the one-step trajectory balance (TB) objective [16, 17]. We use a replay buffer to store the molecular trajectories, which are later used for off-policy training of the generator. Furthermore, we also design a new reward function to provide dense training signals and evaluate molecular trajectories with various lengths between the meta-stable states.
@@
We extensively evaluate our method for three peptides, i.e., Alenine Dipeptide, Polyproline, and Chignolin. We demonstrate that our method faithfully generates a realistic transition path between the given meta-stable states in Figure 1, outperforming the prior ML-based approach [14]. We also show that our method discovers diverse transition states.
@@
Figure 1: Visualization of transition paths generated by our TPS-GFN, progressing from left to right over time. (Top) Polyproline transformation from the meta-stable region PP-II to PP-I. (Bottom) Chignolin folding process.
@@ 2.1 Collective variable-based transition path sampling
Bias potential methods. Bias potential (or force) was originally introduced in enhanced sampling to explore molecular conformations that are difficult to access by the unbiased molecular dynamics (MD) within limited simulation times [18]. We focus on adaptive bias simulations such as meta-dynamics [9] and adaptive biasing force (ABF) [10], which updates the bias potential (or force) adaptively. Meta-dynamics and ABF add bias potential (or force) in a time-dependent manner and enhance the sampling along a few well-designed collective variables (CVs). Other methods for sampling transition paths using bias potentials (or forces) include the umbrella sampling [8], and the Wang-Landau method [19].
MCMC-based methods. Shooting is a well-established method using the Markov chain Monte Carlo (MCMC) procedure on path space [20]. To be specific, it perturbs a shooting point in the transition path, performs MD simulations from it, and proposes new paths. Borrero et al. [21] perturbed the shooting point with meta-dynamics to avoid traps in path space, and Jung et al. [22] trained the parameterized commitor function to select the shooting point with a high acceptance rate. Plainer et al. [23] used latent proposals to leverage information from the latent space of the Boltzmann generator. However, the MCMC-based method suffers from long mixing times to sample independent transition paths, and there is a trade-off between the acceptance rate and diversity. Falkner et al. [24] generated independent shooting points from the Boltzmann generator using the bias potential with the umbrella sampling, but this method relies on CVs and shooting region.
Automated learning of CV. Since many algorithms rely on CVs which require expensive domain knowledge about transitions, researchers have explored automating the derivation of CV from data [25, 26, 27, 28]. Bonati et al. [25] and Trizio et al. [26] used time-lagged independent component analysis and applied deep linear discriminant analysis on MD trajectories. Ray et al. [27] and Yang et al. [28] considered the transition path ensemble as effective data augmentation strategies for discriminant analysis to discover CVs, respectively.
@@ 2.2 Collective variable-free transition path sampling
Generation from a fixed dataset. Given a dataset of transition paths, one can train generative models using the dataset to sample a transition path. Petersen et al. [13] and Lelièvre et al. [12] applied diffusion probabilistic models and [29] variational auto-encoders [30], respectively. However, these methods are limited to small systems, as collecting the transition paths is notoriously difficult due to high energy barriers.
Bias potential-based method. Without any previously collected dataset of transition paths, Das et al. [11] and Hua et al. [31] considered transition path sampling as a reinforcement learning (RL) problem and trained the bias potentials (or forces) of the policy by minimizing KL divergence. However, these works are limited to two-dimensional systems. Notably, Holdijk et al. [14] applied the path integral cross-entropy [32] used in stochastic optimal control framework to scale up to larger systems.
@@ 2.3 Generative flow networks
Generative flow networks (GFlowNets) are the learning framework for amortized inference, generating objects through a series of decisions with a stochastic policy, or forward policy. [15, 16]. By learning the policy to sample an object proportionally to their rewards, GFlowNets can produce diverse samples [15]. To be specific, they train the forward policy by matching its trajectory flow with an auxiliary backward trajectory flow. Training objectives such as flow matching [15], trajectory balance [17], and detailed balance [16] have been proposed. Recent studies have successfully applied GFlowNets to tasks including molecular conformer generation [33], drug discovery [34], crystal structure design [35], and biological sequence design [36].
Comparison with RL and variational inference. RL and variational inference algorithms share similarities with GFlowNets in learning a target distribution specified by a reward function. GFlowNets and RL both make sequential decisions to achieve high rewards, but GFLowNets differ by sampling proportionally to the reward rather than maximizing it. Moreover, while variational inference algorithms aim to learn the target (Boltzmann) distribution, GFlowNets reduce gradient variance by estimating marginal quantities and allow off-policy learning without the need for reweighted importance sampling [37].
@@ 3 Method
Figure 2: The potential energy landscape of Alanine Dipeptide with its collective variable (CV), i.e., two dihedral angles $(\phi,\psi)$. We aim to generate a transition path from one meta-stable state to another, highlighted with yellow arrows.
@@ 3.1 Problem setup
At a high level, our goal is to sample transition paths from one meta-stable state to another meta-stable state through unbiased molecular dynamics (MD), e.g., transition paths from the meta-stable region $C5$ to $C7_{ax}$ in Alanine Dipeptide, as illustrated in Figure 2. However, since such events occur rarely in naïve MD simulation. To this end, we aim to train a neural network to search for such transition paths.
@@
Molecular dynamics. We consider MD simulations that describe motion of a molecular state $\bm{x}=(\bm{r},\bm{v})$ with $N$ atoms, given the atom-wise positions (or conformation) $\bm{r}\in\mathbb{R}^{N\times 3}$ and the atom-wise velocities $\bm{v}\in\mathbb{R}^{N\times 3}$. In particular, we consider Langevin dynamics [38] to describe the molecular motion:
$\mathrm{d}\bm{r}=\bm{v}\mathrm{d}t,\qquad\mathrm{d}\bm{v}=\frac{-\nabla_{\bm{r}}U(\bm{r})}{\bm{m}}\mathrm{d}t-\gamma\bm{v}\mathrm{d}t+\sqrt{\frac{2\gamma k_{B}\lambda}{\bm{m}}}\mathrm{d}\bm{w},$
@@
where $U(\bm{r})$, $\bm{m}$, $\gamma$, $k_{B}$, $\lambda$, and $\mathrm{d}\bm{w}$ denote the potential energy, atom-wise masses, the friction term, Boltzmann constant, temperature, and the Brownian motion, respectively.
@@
Euler Maruyama discretization. Langevin dynamics can be discretized into a sequence of molecular states $\bm{x}_{0},\bm{x}_{1},\ldots,\bm{x}_{T}$ using the Euler Maruyama method [39] as follows:
$\bm{r}_{t}=\bm{r}_{t-1}+\bm{v}_{t-1}\Delta,\qquad\bm{v}_{t}=(1-\gamma\Delta)\bm{v}_{t-1}-\frac{\nabla_{\bm{r}}U(\bm{r}_{t})}{\bm{m}}\Delta+\sqrt{\frac{2\gamma k_{B}\lambda}{\bm{m}}}\bm{\varepsilon}_{t}\Delta,$
@@
where $\Delta$ is the discretization step size and $\bm{\varepsilon}_{t}\in\mathbb{R}^{N\times 3}$ is a noise from the standard normal distribution. We denote the conditional probabilities described by this discretization as $p_{\text{MD}}(\bm{x}_{t}|\bm{x}_{t-1})$.
@@
Transition path sampling. One of the challenges in sampling transition paths through unbiased MD simulations is the meta-stability: a state can remain trapped for a long time in the initial meta-stable region $\mathcal{A}\subseteq\mathbb{R}^{N\times 3}$ before transitioning into a distinct meta-stable region $\mathcal{B}\subseteq\mathbb{R}^{N\times 3}$. Our goal is to sample a transition path $\bm{x}_{0:T}=(\bm{x}_{0},\ldots\bm{x}_{T})$ that links the distinct regions $\mathcal{A}$ and $\mathcal{B}$ where $\bm{r}_{0}\in\mathcal{A}$, $\bm{r}_{T}\in\mathcal{B}$ and $T$ is the length of the trajectory.
@@
Since the meta-stable region $\mathcal{A}$ and $\mathcal{B}$ are not well-specified for many molecular systems, we find unique local minima $\bm{r}_{\mathcal{A}}$ and $\bm{r}_{\mathcal{B}}$ of the potential function inside the meta-stable regions $\mathcal{A}$ and $\mathcal{B}$, respectively. Afterwards, we sample a transition path $\bm{x}_{0:T}$ that starts from the state $\bm{x}_{0}=\bm{x}_{\mathcal{A}}$ and ends nearby the conformation $\bm{r}_{\mathcal{B}}$, i.e., $\bm{r}_{T}\in\mathcal{B}$. To be specific, we fix $\bm{x}_{0}=\bm{x}_{\mathcal{A}}$ and sample the remaining trajectory $\bm{x}_{1:T}$ from the following distribution:
$p_{\mathcal{A},\mathcal{B}}(\bm{x}_{1:T}|\bm{x}_{0})=\frac{1}{Z}\mathbf{1}_{\mathcal{B}}(\bm{r}_{T})\prod_{t=1}^{T}p_{\text{MD}}(\bm{x}_{t}|\bm{x}_{t-1})$ (1)
@@
where $Z$ is the normalizing constant and $\mathbf{1}_{\mathcal{B}}(\bm{r}_{T})$ is the binary indicator function for the final conformation $\bm{r}_{T}$ being in the meta-stable region $\mathcal{B}$. In other words, $\mathbf{1}_{\mathcal{B}}(\bm{r}_{T})=1$ if $\bm{r}_{T}\in\mathcal{B}$, and $\mathbf{1}_{\mathcal{B}}(\bm{r}_{T})=0$ otherwise.

### v3

@@ Abstract
Understanding transition paths between meta-stable states in molecular systems is fundamental for material design and drug discovery. However, sampling these paths via unbiased molecular dynamics simulations is computationally prohibitive due to the high energy barriers between the meta-stable states. Recent machine learning approaches are often restricted to simple systems or rely on collective variables (CVs) extracted from expensive domain knowledge. In this work, we propose to leverage generative flow networks (GFlowNets) to sample transition paths without relying on CVs. We reformulate the problem as amortized energy-based sampling over transition paths and train a neural bias potential by minimizing the squared log-ratio between the target distribution and the generator, derived from the flow matching objective of GFlowNets. Our evaluation on three proteins (Alanine Dipeptide, Polyproline Helix, and Chignolin) demonstrates that our approach, called TPS-GFN, generates more realistic and diverse transition paths than the previous CV-free machine learning approach.
Machine Learning, ICML
@@ 1 Introduction
In material design and drug discovery, understanding the mechanisms and kinetics of transitions between meta-stable states of molecular systems, such as protein folding and chemical reactions (Spotte-Smith et al., 2022; Ahn et al., 2019; Mulholland, 2005; Piana et al., 2012), is crucial. A comprehensive study of them requires sampling diverse transition paths (Lee et al., 2017; Elber, 2016), which provide insights into mechanisms and energy landscapes. However, sampling transition paths through unbiased molecular dynamics (MD) simulations is challenging due to the high energy barriers of intermediate states, which cause an exponential decay in transition probability (Pechukas, 1981). Thus, a brute-force search through unbiased MD simulations is impractical within a realistic computational budget.
@@
To address this problem, researchers have developed bias potential enhanced sampling (BPES) methods such as umbrella sampling (Torrie & Valleau, 1977; Kästner, 2011), meta-dynamics (Ensing et al., 2006; Branduardi et al., 2012; Bussi & Branduardi, 2015), on-the-fly probability-enhanced sampling (Invernizzi & Parrinello, 2020, OPES), and adaptive biasing force (Comer et al., 2015, ABF) methods. BPES methods rely on bias potentials (or forces) to facilitate transitions across high energy barriers. These methods design bias potentials using collective variables (CVs), functions of atomic coordinates that capture the slow modes of the transition. While effective for some systems, the reliance on expensive domain knowledge limits the applicability of BPES methods to systems where CVs are less understood.
@@
Recently, machine learning has emerged as a promising paradigm for both designing and training neural bias potentials—neural network predicting the bias potential—without relying on CVs (Das et al., 2021; Lelièvre et al., 2023; Petersen et al., 2023; Holdijk et al., 2024). For instance, Das et al. (2021) and Lelièvre et al. (2023) used reinforcement learning to sample transition paths from biased molecular dynamics with a parametrized force that approximates the transition path distribution. However, these works are limited to two-dimensional systems. Holdijk et al. (2024) trained the neural bias potential by minimizing KL divergence on the paths collected by itself, i.e., on-policy training. Still, this approach suffers from mode-seeking behavior, and the on-policy training does not reuse the collected paths, leading to inefficiency in sample complexity.
@@
Contribution. In this work, we propose TPS-GFN, generative flow networks (Bengio et al., 2021, GFlowNets or GFNs) for transition path sampling (TPS), which trains the neural bias potential without requiring collective variables (CVs). To this end, we formulate the problem as the energy-based sampling over transition paths between the given meta-stable states. GFlowNet allows off-policy training and avoids collapsing to a particular mode of distribution over transition paths.
@@
In particular, our method trains the neural bias potential by minimizing the squared log-ratio between the target distribution and the generator, which we derive from the one-step trajectory balance (TB) objective (Bengio et al., 2023; Malkin et al., 2022a). We use a replay buffer to store the paths, which are later used for off-policy training of the generator. Furthermore, we also design a new reward function to provide dense training signals and evaluate paths with various lengths between the meta-stable states.
@@
We extensively evaluate our method for three peptides, i.e., Alenine Dipeptide, Polyproline Helix, and Chignolin. We demonstrate that our method faithfully generates a realistic transition path between the given meta-stable states in Figure 1, outperforming the prior ML-based approach (Holdijk et al., 2024). We also show that our method generates diverse transition paths and even discovers transition states.
@@
Figure 1: Transition paths generated by TPS-GFN, progressing from left to right over time. (Top) A conformation change of Alanine Dipeptide. (Middle) An isomerization of Polyproline Helix from left-handed to right-handed helix. (Bottom) A Chignolin folding process.
@@ 2.1 Collective variable-based transition path sampling
Bias potential-based methods. Bias potential (or force) was originally introduced in enhanced sampling to explore molecular conformations that are difficult to access by the unbiased molecular dynamics (MD) within limited simulation times (Hénin et al., 2022). We focus on adaptive bias simulations such as meta-dynamics (Branduardi et al., 2012), on-the-fly probability-enhanced sampling (Invernizzi & Parrinello, 2020, OPES), and adaptive biasing force (Comer et al., 2015, ABF). These methods bias potential (or force) in a time-dependent manner and enhance the sampling along a few well-designed collective variables (CVs). However, selecting collective variables of a molecule requires domain expertise, restricting the applications to systems where CVs are little known.
MCMC-based methods. Shooting is a well-established method using the Markov chain Monte Carlo (MCMC) procedure on path space (Dellago et al., 1998). This method perturbs the velocity of a shooting point—one of the points on a transition path—to initiate MD simulations. It then performs MD simulations from this perturbed point and proposes new transition paths. Borrero & Dellago (2016) moved the shooting point with meta-dynamics to avoid traps in path space. Jung et al. (2023) used a neural network to model commitor function, selecting a shooting point with a high acceptance rate. Plainer et al. (2023) used latent proposals to leverage information from the latent space of the Boltzmann generator (Noé et al., 2019). However, the MCMC-based method suffers from a long mixing time to sample independent transition paths, and there is a trade-off between the acceptance rate and diversity. To avoid long mixing time, Falkner et al. (2022) directly generated independent shooting points from the Boltzmann generator with the umbrella sampling, but this method requires predefined CVs and shooting regions.
Automated learning of CV. To avoid the requirement of expensive domain knowledge about CVs, researchers have explored automating the derivation of CV from data (Bonati et al., 2020, 2021; Trizio & Parrinello, 2021; Ray et al., 2023; Yang et al., 2024). Bonati et al. (2020, 2021) and Trizio & Parrinello (2021) used discriminant analysis and time-lagged independent component analysis, respectively, on the neural network outputs from the physical descriptors collected by MD simulations. Ray et al. (2023) and Yang et al. (2024) considered the transition state ensemble as data augmentation strategies for discriminant analysis.
@@ 2.2 Collective variable-free transition path sampling
Generation from a fixed dataset. Given a dataset of transition paths, one can train generative models using the dataset to sample a transition path. Petersen et al. (2023) and Lelièvre et al. (2023) applied diffusion probabilistic models (Ho et al., 2020) and variational auto-encoders (Kingma & Welling, 2013), respectively. However, these methods are limited to small systems, as collecting the transition paths is notoriously difficult due to high energy barriers.
Neural bias potential-based method. Without a previously collected dataset, Das et al. (2021) and Hua et al. (2024) considered transition path sampling as a reinforcement learning (RL) problem and trained the neural bias potentials (or forces) of the policy by minimizing KL divergence. However, these works are limited to two-dimensional systems. Notably, Holdijk et al. (2024) scaled up to larger systems such as Chignolin and applied the path integral cross-entropy (Kappen & Ruiz, 2016) used in stochastic optimal control framework to train the neural bias potential.
@@ 2.3 Generative flow networks
Generative flow networks (GFlowNets) are the learning framework for amortized inference, generating objects through a series of decisions with a stochastic policy, or forward policy (Bengio et al., 2021, 2023). By learning the policy to sample an object proportionally to their rewards, GFlowNets can produce diverse samples (Bengio et al., 2021). Specifically, they train the forward policy by matching its trajectory flow with an auxiliary backward trajectory flow. Training objectives such as flow matching (Bengio et al., 2021), trajectory balance (Malkin et al., 2022a), and detailed balance (Bengio et al., 2023) have been proposed. Recent studies have successfully applied GFlowNets to tasks including molecular conformer generation (Volokhova et al., 2023), drug discovery (Shen et al., 2023), crystal structure design (Hernandez-Garcia et al., 2023), and biological sequence design (Jain et al., 2022).
Comparison with RL and variational inference. RL and variational inference are similar to GFlowNets in learning a target distribution specified by a reward function. GFlowNets and RL make sequential decisions to achieve high rewards, but GFLowNets differ by sampling proportionally to the reward rather than maximizing it. Moreover, while variational inference aim to learn the target (Boltzmann) distribution, GFlowNets reduce the variance of gradient estimator by estimating marginal quantities and allow off-policy learning without the need for importance sampling (Malkin et al., 2022b).
@@ 3.1 Problem setup
At a high level, our goal is to sample transition paths from one meta-stable state to another meta-stable state through unbiased molecular dynamics (MD), e.g., transition paths from the meta-stable region $C5$ to $C7_{ax}$ of Alanine Dipeptide, as illustrated in Figure 2. Since such events rarely occur due to the high energy barrier between the two meta-stable regions, we train a neural network to search for such transition paths.
@@
Molecular dynamics. We consider MD simulations that describe motion of a molecular state $\bm{x}=(\bm{r},\bm{v})$ with $N$ atoms, given the atom-wise positions (or conformation) $\bm{r}\in\mathbb{R}^{N\times 3}$ and the atom-wise velocities $\bm{v}\in\mathbb{R}^{N\times 3}$. In particular, we consider Langevin dynamics (Bussi & Parrinello, 2007) to describe the molecular motion:
$\mathrm{d}\bm{r}=\bm{v}\mathrm{d}t,$
$\mathrm{d}\bm{v}=\frac{-\nabla_{\bm{r}}U(\bm{r})}{\bm{m}}\mathrm{d}t-\gamma\bm{v}\mathrm{d}t+\sqrt{\frac{2\gamma k_{B}\lambda}{\bm{m}}}\mathrm{d}\bm{w},$
@@
where $U(\bm{r})$, $\bm{m}$, $\gamma$, $k_{B}$, $\lambda$, and $\mathrm{d}\bm{w}$ denote the potential energy, the atom-wise masses, the friction term, the Boltzmann constant, the temperature, and the Brownian motion, respectively.
@@
Euler Maruyama discretization. Langevin dynamics can be discretized into a sequence of molecular states $\bm{x}_{0},\bm{x}_{1},\ldots,\bm{x}_{T}$ using the Euler Maruyama method (Kloeden et al., 2012) as follows:
$\bm{r}_{t}=\bm{r}_{t-1}+\bm{v}_{t-1}\Delta,$
$\bm{v}_{t}=(1-\gamma\Delta)\bm{v}_{t-1}+\frac{-\nabla_{\bm{r}}U(\bm{r}_{t})}{\bm{m}}\Delta+\sqrt{\frac{2\gamma k_{B}\lambda}{\bm{m}}}\bm{\varepsilon}_{t}\Delta,$
@@
where $\Delta$ is the discretization step size and $\bm{\varepsilon}_{t}\in\mathbb{R}^{N\times 3}$ is a noise from the standard normal distribution. We denote the conditional probabilities described above as $p_{\text{MD}}(\bm{x}_{t}|\bm{x}_{t-1})$.
Figure 2: A transition path from the meta-stable state $C5$ to $C7_{ax}$ on the potential energy landscape of Alanine Dipeptide, represented by its well-known collective variables (CV), which are the two backbone dihedral angles $(\phi,\psi)$. The transition path is highlighted with yellow arrows.
@@
Transition path sampling. One of the challenges in sampling transition paths through unbiased MD simulations is the meta-stability: a state remains trapped for a long time in the initial meta-stable region $\mathcal{A}\subseteq\mathbb{R}^{N\times 3}$ before transitioning into a distinct meta-stable region $\mathcal{B}\subseteq\mathbb{R}^{N\times 3}$. Our goal is to sample a transition path $\bm{x}_{0:T}=(\bm{x}_{0},\ldots\bm{x}_{T})$ that links the distinct regions $\mathcal{A}$ and $\mathcal{B}$ where $\bm{r}_{0}\in\mathcal{A}$, $\bm{r}_{T}\in\mathcal{B}$ and $T$ is the length of the path.
@@
Since the meta-stable region $\mathcal{A}$ and $\mathcal{B}$ are not well-specified for many molecular systems, we find unique local minima $\bm{r}_{\mathcal{A}}$ and $\bm{r}_{\mathcal{B}}$ of the potential function inside the meta-stable regions $\mathcal{A}$ and $\mathcal{B}$, respectively. Afterwards, we sample a transition path $\bm{x}_{0:T}$ that starts from the state $\bm{x}_{0}=\bm{x}_{\mathcal{A}}$ and ends nearby the conformation $\bm{r}_{\mathcal{B}}$, i.e., $\bm{r}_{T}\in\mathcal{B}$. To be specific, we fix $\bm{x}_{0}=\bm{x}_{\mathcal{A}}$ and sample the path $\bm{x}_{1:T}$ from the following distribution:
$p_{\mathcal{A},\mathcal{B}}(\bm{x}_{1:T}|\bm{x}_{0})=\frac{1}{Z}\mathbf{1}_{\mathcal{B}}(\bm{r}_{T})\prod_{t=1}^{T}p_{\text{MD}}(\bm{x}_{t}|\bm{x}_{t-1}),$ (1)
@@
where $Z$ is the normalizing constant and $\mathbf{1}_{\mathcal{B}}(\bm{r}_{T})$ is the binary indicator function for the conformation $\bm{r}_{T}$ being in the meta-stable region $\mathcal{B}$. In other words, $\mathbf{1}_{\mathcal{B}}(\bm{r}_{T})=1$ if $\bm{r}_{T}\in\mathcal{B}$, and $\mathbf{1}_{\mathcal{B}}(\bm{r}_{T})=0$ otherwise.
@@
Table 1: Benchmark scores on Alanine Dipeptide which has well-known CVs. All metrics are averaged over 64 trajectories. ETP and EFP are reported only if the final state is in the target meta-stable state. The best values are highlighted in bold except unbiased MD (300K). Values from the paper Holdijk et al. (2024) and the model parameter of its official code are indicated by $\dagger$ and $\ddagger$, respectively. Methods using bias potential and force are denoted by (P) and (F), respectively.
[…]
MD (300K) $7.52\pm 0.96$ $0.00$--- 00.5.72 $\pm$ 0.00 $84.73\pm 4.97$-
MD (9000K) $7.96\pm 2.36$ 04.684.684.684.68 --$870.11\pm 164.52$ $-13.14\pm 2.83$ 0115.06±31.37plus-or-minus115.0631.37115.06$\pm$ 31.37115.06 ± 31.37 $789.54\pm 134.34$
PIPS (P) $1.21\pm 0.31$ $63.50$ 00--8.35 $\pm$ 8.0400 00.-- 0-- --
[…]
TPS-GFN (F) $1.48\pm 0.21$ $81.25$ 0.$11.66\pm 18.44$ 0--$5.57\pm 0.02$ 08.77±4.12plus-or-minus8.774.128.77$\pm$ 4.128.77 ± 4.12 30.14±11.46plus-or-minus30.1411.46-30.14$\pm$ 11.46- 30.14 ± 11.460.
Table 2: Benchmark scores on Polyproline Helix and Chignolin. All metrics are averaged over 64 trajectories. The best values are highlighted in bold except unbiased MD. Reproduced results from Holdijk et al. (2024) are indicated by $\dagger$. Methods using bias potential and force are denoted by (P) and (F), respectively.
Method EPD ($\downarrow$) LL ($\uparrow$) EPCD ($\downarrow$)
$\mathrm{nm\times 10^{-3}}$$\times 10^{2}$ $\mathrm{nm}$
Polyproline Helix
MD (300K) $24.14\pm 3.10$ $13.89\pm 0.01$ $87.96\pm 3.93$
PIPS (P) $14.81\pm 5.32$ $13.77\pm 0.02$ 073.83±17.94plus-or-minus73.8317.9473.83$\pm$ 17.9473.83 ± 17.94
PIPS (F) $30.25\pm 3.29$ $13.75\pm 0.02$ $104.89\pm 4.60$
[…]
TPS-GFN (F) $17.63\pm 1.32$ $\textbf{13.87}\pm\textbf{0.02}$ $82.08\pm 4.02$
Chignolin
MD (300K) $367.84\pm 101.43$ $36.01\pm 0.04$ $258.49\pm 21.15$
PIPS (P) $59.73\pm 11.19$ $34.03\pm 0.44$ $184.97\pm 18.57$
PIPS (F) $90.60\pm 43.86$ $33.83\pm 0.33$ $214.42\pm 39.63$
[…]
TPS-GFN (F) $55.21\pm 10.15$ $\textbf{34.54}\pm\textbf{0.14}$ $148.34\pm 18.14$